A harmonic oscillator has a mass of 0•20kg,an unknown spring constant, if the period is T=20. Find k


1
Expert's answer
2022-03-21T18:49:39-0400

"\omega=\frac{2\pi}T=0.31~\frac{rad}s,"

"k=\omega^2 m=0.02~\frac N{kg}."